WAVEGUIDE CHIRPED-PULSE FOURIER TRANSFORM MICROWAVE (CP-FTMW) SPECTRUM OF ORTHO-FLUOROTOLUENE

IAN A. FINNERAN and \underline{STEVEN T. SHIPMAN}, Division of Natural Sciences, New College of Florida, Sarasota, FL 34243.

The microwave spectrum of o-fluorotoluene has been measured at 0 ^{∘}C from 8.7–18.3 GHz with waveguide chirped-pulse Fourier transform microwave spectroscopy (CP-FTMW). We have extended previous assignments of the lowest energy Aand E-states by Susskind^{a} and Mäder^{b} and report on preliminary assignments of vibrationally excited states. This molecule also serves as a proof-of-principle for “coarse” microwave-microwave double resonance (MW-MW DR) measurements, in which we pump all transitions within a relatively broad frequency range (> 50 MHz) simultaneously. Transitions connected to peaks within the bandwidth of the coarse pulse are intensity modulated, revealing approximate connectivities that can be refined by performing “fine” MW-MW DR measurements. Prospects for using this method to perform automated, time-efficient MW-MW DR on samples with dense spectra and unknown assignments will be discussed.

aSusskind, J., J. Chem. Phys. 53, 2492 (1970).

bJacobsen, S., Andresen, U., and Mäder, H., Struct. Chem. 14, 217 (2003).